If n(A) = 22, n(B) = 13, nd n(A ∩B) = 6 then n(A U B)=
a. 35
b. 28
c. 29
d. 27
: If n(A) = 22, n(B) = 13, nd n(A ∩B) = 6 then n(A U B)=
a. 35
b. 28
c. 29
d. 27

Answer by stanbon(19680)

If n(A) = 22, n(B) = 13, and n(A and B) = 6 then n(A U B)=
a. 35
b. 28
c. 29
d. 27
----------------
n(AUB) = n(A)+n(B)-n(A and B)
n(AUB) = 22 + 13 - 6 = 29
======================
Cheers,
Stan H.